OpenStudy (anonymous):

If this calculation continued forever, what would you expect the answer to be: 1/3 + 1/9 + 1/27 + 1/81 + 1/243 + ...

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{1}{3^2}+\frac{1}{3^3}+\frac{1}{3^4}+...\] so you know how to add this up ?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

\[a+ar+ar^2+ar^3+ar^4+...=\frac{a}{1-r}\] use \[a=r=\frac{1}{3}\]

OpenStudy (anonymous):

you can almost do it in your head almost what is one minus one third?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

ok good what is the reciprocal of \(\frac{2}{3}\) ?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

and finally, what is \(\frac{3}{2}\times \frac{1}{3}\) ?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

im trying to understand the formula how it works

OpenStudy (anonymous):

the formula i wrote? \[\frac{a}{1-r}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you mean what to plug in for \(a\) and \(r\)? or do you mean why the formula works? or do you mean how you compute using the formula?

OpenStudy (anonymous):

what is a and what is r?

OpenStudy (anonymous):

\(a\) is the first term, and \(r\) is what you multiply the first term by to get the second term what you multiply the second term by to get the third term, and so on

OpenStudy (anonymous):

this formula only works if you have a sum that looks like \[a+ar+ar^2+ar^3+ar^4+...\]

OpenStudy (anonymous):

in your example you do you have \[\frac{1}{3}+\frac{1}{3}\times \frac{1}{3}+\frac{1}{3}\times \left(\frac{1}{3}\right)^2+...\]

OpenStudy (anonymous):

in other words, you start with \(\frac{1}{3}\) as the first term and keep multiplying each successive term by \(\frac{1}{3}\) it is not always the case that \(a=r\) but it is in your example

OpenStudy (anonymous):

perfectly explained thanks alot @satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

on e last question: how did you deduct the formula?

OpenStudy (anonymous):

what if it is 1/3 + 1/9 - 1/27 + 1/81 - 1/243 + ...

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

drop off the first \(\frac{1}{3}\) and it looks like \(a=\frac{1}{9}\) and \(r=-\frac{1}{3}\)

OpenStudy (anonymous):

compute \[\frac{\frac{1}{9}}{1+\frac{1}{3}}\] and then add \(\frac{1}{3}\) to the result

OpenStudy (anonymous):

could you please explain the formula?

OpenStudy (anonymous):

why is a =1/9

OpenStudy (anonymous):

1 - 1/2 + 1/4 - 1/8 + 1/16 - 1/32 + 1/64 - ...

OpenStudy (anonymous):

this is the problem i cant get

OpenStudy (anonymous):

you still there? i was off on a tangent

OpenStudy (anonymous):

\[1 - 1/2 + 1/4 - 1/8 + 1/16 - 1/32 + 1/64... \] it alternates so \(r\) is negative

OpenStudy (anonymous):

\(a=1,r=-\frac{1}{2}\)

OpenStudy (anonymous):

compute as before \[\frac{a}{1-r}\] with \(a=1,r=-\frac{1}{2}\) to get \(\frac{1}{1+\frac{1}{2}}\)

OpenStudy (anonymous):

why is a=1 and r=-1/2?

OpenStudy (anonymous):

but does the same logic apply for the first question @satellite73

OpenStudy (anonymous):

1/3 + 1/9 + 1/27 + 1/81 + 1/243 + ...

OpenStudy (anonymous):

will 1/3 be the first term (a=1/3)???

OpenStudy (anonymous):

what will r be? 1st term multiply 2nd term (1/3 x 1/9)

OpenStudy (anonymous):

@satellite73